Use the method of variation of parameters to determine the general solution of the given differential equation.

y′′′−y′=5t

Use C1, C2, C3, ... for the constants of integration.

y(t) = _______
1
Expert's answer
2020-07-07T19:28:49-0400

The given differential equation can be written as "(D^{3}-D)y = 5t,~ \text{where}~ D\equiv \dfrac{d}{dt}."


The auxiliary equation is, "m^{3}-m=0." Solving, we get "m = 0,1,-1."


The complementary function is, "y_{c} = c_{1}+c_{2}e^{t}+c_{3}e^{-t} = c_{1}f_{1}+c_{2}f_{2}+c_{3}f_{3}"


The particular integral is "y_{p}=Pf_{1}+Qf_{2}+Rf_{3}," where


"P=\displaystyle\int \frac{W_{1}}{W(f_{1},f_{2},f_{3})}dx; Q=\displaystyle\int \frac{W_{2}}{W(f_{1},f_{2},f_{3})}dx;\\~\\ R=\displaystyle\int \frac{W_{3}}{W(f_{1},f_{2},f_{3})}dx;"


and "W_{i}" is the determinant obtained from the Wronskian "W(f_{1},f_{2},f_{3})" by replacing the "i^{th}" - column by the vector "\begin{pmatrix} 0 \\ 0 \\ g(t) \end{pmatrix}" where "g(t) = 5t."

Now,


"W(f_{1},f_{2},f_{3}) = \begin{vmatrix} 1 & e^{t} & e^{- t} \\ 0 & e^{t} &- e^{- t} \\0 & e^{t} & e^{- t} \end{vmatrix} = 2\\~\\
W_{1} = \begin{vmatrix} 0 & e^{t} & e^{- t} \\ 0 & e^{t} &- e^{- t} \\5t & e^{t} & e^{- t} \end{vmatrix} = -10t\\~\\
W_{2} = \begin{vmatrix} 1 & 0 & e^{- t} \\ 0 & 0 &- e^{- t} \\0 & 5t & e^{- t} \end{vmatrix} = 5te^{-t}\\~\\
W_{3} = \begin{vmatrix} 1 & e^{t} & 0 \\ 0 & e^{t} & 0 \\0 & e^{t} &5t \end{vmatrix} = 5te^{t}"


Therefore,


"P = \displaystyle\int -5t~ dt = -\frac{5t^{2}}{2}\\~\\
Q= \displaystyle \frac{5}{2}\int te^{-t}~ dt = \frac{5}{2}(-te^{-t}-e^{-t}) = -\frac{5}{2}e^{-t}(t+1)\\~\\
R= \displaystyle \frac{5}{2}\int te^{t}~ dt = \frac{5}{2}(te^{t}-e^{t}) = \frac{5}{2}e^{t}(t-1)"


Hence,


"\begin{aligned}
y_{p} &=Pf_{1}+Qf_{2}+Rf_{3}\\
&=-\frac{5t^{2}}{2}-\frac{5}{2}e^{-t}(t+1) \cdot e^{t} + \frac{5}{2}e^{t}(t-1) \cdot e^{-t}\\
&=-\frac{5t^{2}}{2}-\frac{5}{2}(t+1) + \frac{5}{2}(t-1)\\
&=-\frac{5t^{2}}{2} - 5
\end{aligned}"


The general solution is

"\begin{aligned}
y = y_{c}+y_{p} &= c_{1}+c_{2}e^{t}+c_{3}e^{-t}-\dfrac{5t^{2}}{2}-5\\
y&= C_{1}+c_{2}e^{t}+c_{3}e^{-t}-\dfrac{5t^{2}}{2}, \text{where}~C_{1}=c_{1}-5.
\end{aligned}"